Pb(NO3)2 + 2 NaI à PbI2 + 2 NaNO 

7)   How many grams of the product were formed in no. 6 if the percent yield is 80%?

1
Expert's answer
2022-03-08T16:54:04-0500

Pb(NO3)2 + 2 NaI = PbI2 + 2 NaNO

If we had 5 g of Pb(NO3)2

5 g Pb(NO3)2 - x g PbI2

331.2 g/mol Pb(NO3)2 - 461.01g/mol PbI2

x = 461.01 × 5 / 331.2 = 6.96 g

m(PbI2) = 0.8 × 6.96 g = 5.57 g

5 g Pb(NO3)2 - y g NaNO3

331.2 g/mol Pb(NO3)2 - 2 mol × 84.99 g/mol NaNO3

y = 2 × 84.99 × 5 / 331.2 = 2.57 g

m(NaNO3) = 0.8 × 2.57 g = 2.06 g